OpenStudy (anonymous):

What is the simplified form of the expression 8y - 6y + 2x + y + 9? a. 2x + 15y + 9 b. 2x + 3y + 9 c. 5xy + 9 d. 14xy

OpenStudy (colouringme):

B

OpenStudy (anonymous):

Thanks :)! btw like your picture haha

OpenStudy (colouringme):

=(8-6+1)y + 2x + 9 =3y + 2x + 9 = 2x + 3y + 9 Hahah,no problem, Darwin is funny. ((: , why don't you put AT related profile pics ? (: